OpenStudy (christianx2):

I'm working on some algebra 2 homework and I got this problem for number 51: Find a1 for each Sn = -2912, r = 3, n = 6. I plugged this into the formula to get 2912 = a1 - a1(792)/2 and then multiplied the 2 to get 5824 = a1-a1(729), what do I do from here?

jimthompson5910 (jim_thompson5910):

Make sure you fix the 792 and change it to 729

jimthompson5910 (jim_thompson5910):

if it helps, think of a1 as x so you'd have 5824 = x - x*729

OpenStudy (christianx2):

But then I would I do 5824/729 and get 7.989 = x - x and then 7.989=0

jimthompson5910 (jim_thompson5910):

what would x - x*729 simplify to?

jimthompson5910 (jim_thompson5910):

hints x = 1x x*729 = 729x

OpenStudy (christianx2):

OOOOhhhhhhh I see it now

OpenStudy (christianx2):

728x?

jimthompson5910 (jim_thompson5910):

small - big = negative number

jimthompson5910 (jim_thompson5910):

1-729 = -728

OpenStudy (christianx2):

Ok so -728x = 5824? then 5824/-728 = -8 = a1 right?

jimthompson5910 (jim_thompson5910):

yep the first term is \(\Large a_1 = {\color{red}{-8}}\)

OpenStudy (christianx2):

Ok I got it now. Thanks for your help Jim.

jimthompson5910 (jim_thompson5910):

you're welcome